2sin(theta)-5+2csc(theta)=0 when 0<(theta)<90 degrees
Answer by stanbon(48558)

2sin(theta)-5+2csc(theta)=0 when 0<(theta)<90 degrees
----
Multiply thru by sin(theta) to get:
2sin^2(theta) - 5sin(theta) + 2 = 0
Factor:
(2sin(theta) -1)(sin(theta)-2) = 0
---
sin(theta) = 1/2 or sin(theta) = 2
-----
Only sin(theta) = 1/2 is possible.
Theta = 30 degrees
==========================
Cheers,
Stan H.